Hurdle Corporation is authorized to issue 200,000 shares of common stock. In its only issuance to...

  1. Hurdle Corporation is authorized to issue 200,000 shares of common stock. In its only issuance to date, Hurdle sold 50,000 shares. A few months later, Hurdle reacquired 5,000 of those issued shares with the intention of reissuing them all later. By the end of this year, Hurdle had reissued just 2,000 of the reacquired shares. The board of directors declared a 40-cent dividend at the end of the year. How many shares will receive a dividend?
    1. 200,000
    2.   50,000
    3.   45,000
    4.   47,000
  2. Ernie Banks owned 50 shares of Cub Corporation $2 par value common stock prior to the company enacting a 2 for 1 stock split. How many shares does Ernie own after the split?
    1. 100
    2.   50
    3.   25
    4. 200
  3. PROBLEM 15 (15 pts.)

    Lambert Corporation has 8,000 shares of 10%, $5 par value preferred stock and 100,000 shares of common stock outstanding. The preferred stock is cumulative. Prepare a table showing the amount of

    Dividends allocated to each group of stockholders each year using the following information.

       YEAR         TOTAL DIVIDENDS

        2017                  $3,000

        2018                  $4,500

       

        2019                  $6,000

    PROBLEM 16   (24 pts.)

    On January 1 of this year, Claypool, Inc.had 50,000 shares of its $1 par value common stock issued and outstanding. Prepare general journal entries for the following transactions involving its stock.

    March 8    Claypool reacquired 7,000 of its own shares with plans to reissue them at a later

         date. Claypool paid $8 per share.

    April 7      Reissued 2000 shares for a total of $15,000.    ($7.50 per share)

    May 11     Reissued 1000 shares for a total of $8,500.   ($8.50 per share)

    June 14     Reissued 2000 shares for a total of $15,800.     ($7.90 per share)
Answer #1

1. Dividend will be paid to outside shareholders only and not for treasury shares. Thus, shares that qualify for dividend are: 50,000 - 5,000 (Reacquired) + 2,000 (Reissued) = 47,000. Hence Option D is Correct.

2. Since the split is 2 for 1 share, Ernie Bank will own 50*2 = 100 shares after stock split. Thus Option A is correct.

3. Preference Dividend is cumulative ie, 5*8000*10% = $ 4,000

10% Preference Share Common Stock
2017 3000 0
2018 1000 (2017)+ 3,500 (2018) = 4500 0
2019 500 (2018) + 4000 = 4,500 6000-4500 = 1,500

4.

Journal Entries($)
March 8 Treasury Stock Dr 56,000
To Cash 56,000
(Shares repurchased)
Cash Dr (7.5*2000) 15,000
Additional Paid in capital excess of par Dr (8-7.5)*2000 1,000
To Treasury Stock 16,000
(Treasury stock Reissued)
May 11 Cash Dr(8.5*1000) 8,500
To Treasury Stock 8,000
To Additional Paid in capital excess of par (8.5-8)*1000 500
(Treasury stock Reissued)
June 14 Cash Dr (7.9*2000) 15,800
Additional Paid in capital excess of par Dr (8-7.9)*2000 200
To Treasury Stock 16,000
(Treasury stock Reissued)